ganeshie8 (ganeshie8):

lens in a camera and 1/p + 1/i = 1/f formula

ganeshie8 (ganeshie8):

1) for a particular camera, the focal length of its lens is fixed. 2) the image location of the light passing through a convex lens depends on the object location by the formula 1/p + 1/i = 1/f

ILovePuppiesLol (ilovepuppieslol):

if i may ask? What level of education is this?

ganeshie8 (ganeshie8):

this is 12th grade physics @ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

oh my goodness, let me just go to bed *slowly backs away*

ILovePuppiesLol (ilovepuppieslol):

doing my 9th grade biology ;~;

ganeshie8 (ganeshie8):

Here is the question : The image location must change based on the distance the object is from the lens. But the location of the image sensor in a camera is fixed; then, how are we able to take pictures of an object at any arbitrary distance before the lens ?

ganeshie8 (ganeshie8):

Nice... this isn't hard @ILovePuppiesLol and I guess this is more of biology too Camera works more or less same as the human eye. your eye is the camera, your retina is the image sensor... Human eye is able to focus object at any depth by changing the focal length of the lens. The muscles attached tot he lens contract and expand as you try to focus on objects based on the distance. But in a camera, the size of the lens is fixed; so the focal length is also fixed. I'm trying to understand how the "focus" in camera works...

ILovePuppiesLol (ilovepuppieslol):

i thought this was some sort to of matrix math that they teach in mathematician school

OpenStudy (kainui):

Is that really how our eyes work? I thought we focused by moving our eyes closer or further apart. But I guess if it's really our iris contracting that focuses then I guess by analogy a camera's aperture serves the same function: https://i.ytimg.com/vi/F8T94sdiNjc/maxresdefault.jpg

ganeshie8 (ganeshie8):

aperture controls the amount of light passing through the lens right ?

ganeshie8 (ganeshie8):

small aperture requires less shutter speed so as to allow longer exposure times

ganeshie8 (ganeshie8):

how does that change the location of image ?

OpenStudy (kainui):

I don't know, I just thought like by analogy they are mechanically the same sorta shape and probably have the same sorta effect. I don't know how it works, just throwing that out there but I'm working on something else right now so I'll take a more serious approach later.

Miracrown (miracrown):

location of image sensor in camera is fixed, but,..the distance between the lens and the image is changing, because you zoon the lens, is it not? are you able to zoom the lens?

Miracrown (miracrown):

if so, the distance between the sensor and the lens is changing, though, the sensor itself is fixed

ganeshie8 (ganeshie8):

1/p + 1/i = 1/f if `i`(image location) is fixed and `p`(object location) varies, then `f`(focal length) of the lens has to change right ?

Miracrown (miracrown):

no, f of lens does not change,

Miracrown (miracrown):

the first part of the sentence is the misleading one

OpenStudy (baru):

@ganeshie8 i'm guessing you are thinking of those old cheap film cameras that dont come with any lens adjustment options.. ?

Miracrown (miracrown):

if i(image location) is fixed,...that is very misleading.

ganeshie8 (ganeshie8):

yeah, so i have to zoom in / out to be able to focus objects at different distances ?

Miracrown (miracrown):

One should ask the following question: how is the distance i measured? from where is it measured? i is distance from the lens to the image location in camera

ganeshie8 (ganeshie8):

it is shown in my attachment...

Miracrown (miracrown):

if Camera is fixed in your hand, image location is fixed in your perspective. But, distance between image plane and lens is not fixed

Miracrown (miracrown):

yes,..if you zoom to get objects from different distances in focus, then, i will change

Miracrown (miracrown):

but, one's intuitive feeling of image being fixed inside your camera conflicts with the formula, and how the distances are measured in the formula

ganeshie8 (ganeshie8):

may be to keep things simpel... lets stick to a lens w/o zoom

Miracrown (miracrown):

so, it is always measured from the lens

ganeshie8 (ganeshie8):

then i can capture only objects at one particular distance is it ?

Miracrown (miracrown):

if its without zoom, then, it can correctly image only one distance

Miracrown (miracrown):

If you have digital processing enabled, that is another story altogether,...

ganeshie8 (ganeshie8):

thanks, that is all i wanted to know..

Miracrown (miracrown):

but, in the absence of that, yes, you can capture only objects at one particular distance

Miracrown (miracrown):

yw neshie

ganeshie8 (ganeshie8):

looks the aperture controls the blurness too as we reduce the aperture, the rays from a point source entering the lens become more parallel, giving a sharper image

ganeshie8 (ganeshie8):

looks i need to take a quick photography course to understand these better